### v2

# Does quantum information require additional structure?

Abstract
We consider interpretative problems of models of physical reality, including quantum information phenomenon in the context of Heisenberg’s classification of the fundamental physical theoretic models under the role of universal constants Planck’s constant $h$ and speed of light $c$. We introduce the correspondence principle between physical reality and mathematical models, and we discuss its significance in relation to physical models. We consider the status of quantum information in the standard quantum model, and based on the correspondence principle, we propose an interpretation of the wave function as a mathematical representation of quantum information. In this context, we consider Clauser’s analysis of incompatibility formulations of quantum theory in laboratory space and configuration space in the context of local realism. Then, we introduce the hypothesis of quantum space of directly unobserved relations, which precede correlations in the “classical” Minkowski space-time and are compatible with the Reichenbach common cause principle. Finally, we present the Chyliński model as an example of quantum relational space, which predicts potentially measurable effects for the bound states.
keywords:
quantum information, wave function, correspondence principle, quantum relational space
“Heisenberg’s motivation for studying physics
was not only to solve particular problems but also to
illuminate the discussion of broad philosophical question”
Frank Wilczek “Four Big Questions with Pretty Good Answers”
1 Introduction
Quantum theory is the best theory we have to describe physical reality. “All our former experience with application of quantum theory seems to say: what is predicted by quantum formalism must occur in laboratory” [1]. It can be said that it works very well, which is why some physicists claim that “Quantum theory needs no interpretation” [2]. Nevertheless, there is a significant group of scientists for whom the lack of a coherent, commonly accepted interpretation is a source of discomfort. There have been many significant attempts to build a consistent interpretation, such as Copenhagen interpretation, Hidden Variable interpretations [3, 4, 5], and Relativistic interpretations [6, 7, 8]. However, there is still no consensus on what quantum mechanics is. There have been various propositions: “only on relative facts” [8], “only on correlations and only correlations” [9], “on the only fields” [10], “All is $\psi$[11], “The everything-is-a-quantum-wave” [12].
One of the obstacles to a deeper understanding of quantum theory is that it radically deviates from classical patterns. Even pioneer quantum mechanics formulated nearly a hundred years ago, it appeared as a kind of quantum inscription encoded in the language of Hilbert space - a set of rules specifying how to make probabilistic predictions of results of future measurements in laboratories. The scale of difficulty in its deciphering reflects the fact that the discoveries of such phenomena as EPR correlations (entanglement), quantum computing [13, 14], quantum teleportation [15], existence the quantum correlations beyond entanglement [17, 16, 18] took place long after its birth. Quite surprasingly despite winning the battle with nature in an experimental demonstration of breaking Bell’s inequality (Nobel Prize in physics 2022) “…there it still no concensus about what such violations actually imply” [19] (and ref. there in). Moreover, quantum theory does not offer any simple footbridge from the quantum world to our actual world (see in this context [20, 21, 22, 23, 24, 25, 26, 27]. This issue covers many aspects: it has a long history and a vast literature, including the famous Wigner paradox [28, 29, 30].
Paradoxically, although the concept of quantum information lies at the heart of quantum theory, it has rather little impact on its interpretative problems. The vast majority of considerations about the meaning of quantum formalism do not take into account the problem of the relationship between physical reality and mathematical models, including the role of universal constants. This work does not aim to solve these drawbacks of quantum theory but rather attempt to consider the quantum inscription in the context proposed here the correspondence principle and the concept of quantum information.
In sec.2, we introduce the correspondence principle between physical reality and mathematical models and we discuss its significance in relation to physical models (theories) in the context of their classification under the roles of universal constants. In sec.3, we briefly discuss the relativistic classical model and the standard (nonrelativistic) quantum model, emphasizing the roles of universal constants and symmetry groups. In sec.4, we consider the status of quantum information in the standard quantum model, and based on the correspondence principle, we propose an interpretation of the wave function as a mathematical representation of quantum information. In sec.5, we consider Clauser’s analysis of incompatibility formulations of quantum theory in laboratory space and configuration space in the context of local realism. In sec.6, we hypothesize that quantum information requires the existence of an intrinsic relational quantum space in which reversible premeasurements take place and which precedes the space of correlations actualized in Minkowski space. In sec.7, we present Chyliński’s model of quantum relational continuum space, which leads to measurable effects and, at the level of operational observation, is consistent with special relativity.
2 Heisenberg’s classification of fundamental models and the correspondence principle
Looking at the impressive development of physical sciences, a fundamental question constantly comes back to us: What is the reason why nature allows itself to be described by means of mathematical structures and why, out of many potential structures, it accepts only certain selected ones, e.g. Riemannian geometry but non-set-theoretic geometry with Banach-Tarski paradox [31]. This peculiar selectivity of nature can be naturally related to the existence of an active observer, equipped with mathematical tools, which organize his rational thinking about the world and allow him to create models of physical reality. By the term active observer, we mean the conscious observer endowed with free will who can have access to the system directly or indirectly via control and measurement interactions associated with an experimental setup. However, the action of an active observer is accompanied by the activity of nature itself due to the existence of fundamental interactions. Hence, the potential connection between the mathematical structure of models and the structure of the described reality is far from obvious [32, 33, 19]. Two approaches can be distinguished [33]: the ”theory-driven” approach, which assumes that the production of experimental data is governed by a pre-existing mathematical model and the ”data-driven” approach according to which experimental data are considered to be immutable facts, which require to construct corresponding mathematical models. In this context, it is tempting to introduce a correspondence principle (CP):
Any consistent successful partial description of nature is a sort of correspondence between the structure of physical reality and the structure of its mathematical representation. Here, by the term “physical reality,” we mean a set of empirical observer-independent data, which includes all fundamental physical constants [34]. The term “partial description of nature” expresses the fact that, in practice, an observer has access to only a part of the world called a “system”. The term “correspondence” means that physical reality expressed in the language of specific mathematical structures is true according to the theory of reflection [35]. The problem is that the correspondence principle itself does not specify how elements of mathematical structure are related to physical reality. Therefore, the observer must somehow establish a set of interpretative statements to extract meaning from mathematical formalism. He can always, in spirit Gödel, supplement the model with additional structure before applying it [36].
It is significant that “canonical” models of reality, including the Standard Model need constants parameters (fundamental constants) that cannot be explained within the framework of the given model, and they can only be determined from the experiment. Note that the correspondence principle allows for an “operational” interpretation of the model-independent parameters, as interaction constants between the mathematical structure of a model and the structure of the reality it describes. These constants impose global realistic constraints on the mathematical structure of models. There are the “conditional quantities” necessary for the model to reproduce the experimental results. In particular, they build relationships between quantities that were thought to be incommensurable.
Note that there is no characterization of fundamental constants, and their list usually depends on what we accept as a fundamental model (theory). Hence, not all constants are on the same footing. Historically, the Planck’s constants $h$, and speed of light $c$ promoted to the role of universal constants. Here, we will use the term “universal constants” for which we assume that they meet the universality postulate: values of universal constants measured in any reference system always take the same values (see in this context[37, 38, 39]).
According to Heisenberg, the introduction of universal constants led to abrupt changes in the paradigms of four fundamental models (theories) [40]: i) the non-relativistic classical (Newtonian model) (NRC): $01/c=$ 0, $h=0$, ii) relativistic classical (RC): $1/c\neq 0$ and $\hbar=0$, iii) the non-relativistic quantum (NRQ): $01/c=$ 0, $\hbar\neq 0$ iv) the relativistic quantum field (RQF): $1/c\neq 0$ and $\hbar\neq 0$. This classification does not take into account the gravitational constant $G$, which may be of fundamental importance in the quantum-to-classical transition [41].
Note that there were also attempts to build a quantum relativistic models with $h/c\neq 0$ [42, 43, 44]. In the final part of the paper, we will present Chyliński’s relation-space-time model $h/c\neq 0$ [43] which, as far as we know, has not been considered so far, in any physical or philosophical context.
To our knowledge, there is no commonly accepted characterization of the basic models (theories). Here, we adopt a simple characterization that defines the underlying model as closed and compact [45]. We say that a given model (theory) is closed if the internal symmetry of its equations is the symmetry of its measurement space-time (empty background), and it is compact if it explains the structure of measures and clocks necessary to create its metric background.
3 Relativistic classical model
The RC based on the Galileo-Einstein principle and the observer-independent of the speed of light $c$ leads to the concept of Minkowskian space-time $M$ and its Poincare symmetry group as a relativistic model of space-time. Note that in the RC the measurement is passive in the sense that it does not disturb the system such that information can be freely copied by repeatedly interacting with the system. It can be explained by assuming that the observer only has at his disposal “classical” information carriers (CIC) which carry an arbitrarily small portion of energy and momentum. Hence, the recoil of the system being “measured” by the CIC is reduced to zero. It makes the RC non-contextual, as the reproducible properties of the system can be defined independently of any context. It should be noted here, however, that if one assumes that an active observer can freely influence physical systems without pre-defined correlations, this always generates contextuality [32].
Note that the mathematical structure of the canonic phase space can be (according to CP principle) directly attributed to the physical structure as the probability distributions assigned to the classical phase space regions unambiguously correspond to the possible states of the system in which it has specific properties. Clearly, the RC is not compact. As Einstein already noted: “in the present state of theoretical physics, measures, and clocks must be considered as independent concepts, and we are far from a precise knowledge of theoretical foundations to be able to provide the structure of these objects.”
4 Non-relativistic (standard) quantum model
In contrast to the classical models, the role of CP in relation to the quantum ones is much more subtle, as that NRQ strongly deviates from classical patterns. It appears rather as a quantum inscription written in the language of Hilbert space where the observables represented by Hermitian’s operators and the quantum states by positive unit trace matrices cannot be directly attributed to the physical reality. This step changes the paradigm in the description of nature has its origins in the fact that the mathematical structure of the NRQ includes Planck’s constant $h\neq 0$, which i) determines the ultimate limit on the non-commutativity of measurement operators that impose constraints on the possibility of extracting information by measurement from a quantum system[46, 47], and separation of the measured system (“specimen”) from the measuring device ceases to be obvious as in the RC. It implies ii) breaking the rescaling of the spatiotemporal extension of the systems by determining the finite volume of the phase cell, iii) the existence of the quantum information carriers (QIC),
The NRQ model applies to the phenomenological regime in which energies are low, and the effects of special relativity and high-energy physics can be neglected. The relativistic effects can be described within the framework of the RQ, which is based on the Lorentz-Poincare symmetry $M_{4}$. Note, however, that the NRQ model is fundamental because it is closed and compact. Closed theories in Galilean $G_{4}$ space-time can introduce hypothetical particles characterized by their absolute masses and spins [48]. In particular, the intrinsic symmetry of the non-relativistic Heisenberg’s, Schrödinger’s, and Pauli’s equations is also the symmetry of their empty Galilean group $G_{4}$ background. As one knows, the unitary representations of the $G_{4}$ have a nontrivial projective nature with physical consequences: Bargmann’s superselection mass rule [49], the noncommutativity of operators momentum and position, the possibility of a spin angular momentum (Pauli’s model). The “elementary” particles are defined by irreducible projective unitary representations of the Galilean group in some Hilbert space, which do not interact directly but via a field that mediates the interaction between particles [50, 49].
It is generally accepted that the NRQ is neither universal nor the best model of reality, but it is certainly fundamental, and it can be a starting point for more advanced models such as RQ or S-matrix formalism. It is compact in the sense that it reproduces the extent of stable macroscopic measures with a hierarchical structure (quarks, nucleons, nuclei, atoms, molecules…). Since NRQ was completed, there have been times when physicists thought that its “quantum code” was deciphered completely, like the Rosetta Stone. Therefore, a big surprise was when Einstein Podolski and Rosen [51] discovered that NRQ predicts the existence of “spooky” correlations between properties of particles regardless of distance called by Schrödinger ”Verschrankung” (entanglement) [52]. Even though more than half a century has passed since its discovery, its full physical meaning still remains far from clear. Recently, it was argued that “…a violation of Bell-type inequalities by experimental data merely reflects the properties of the process, chosen by the experimenter, to select groups of data, not an intrinsic property of the data themselves.”[19]. Quantum entanglement and Born law break fundamental features of the RC: the separability and classical determinism, but the full significance of these facts seems to require new concepts.
Interestingly, the NRQ predicts a number of quantum phenomena without any reference to RQ: electron diffraction in the Davisson-Germer experiment, the quantum Hall effect, scattering experiments which involve collisions where radiative effects can be ignored [36]. However other well-known quantum phenomena, such as the two-slit experiment random decay times and neutron diffraction cannot be described completely within the framework of NRQ without referring to RQ if the detection processes are included.
5 Quantum model as a paradigm for quantum information
There have been many obstacles to deciphering the potential physical implications of rich structure NRQ. It was only a series of groundbreaking theoretical and experimental papers showed that at a fundamental level, nature reveals the existence of a new, subtle property of physical systems - non-classical (or quantum) information [53], on the processing of which the mathematical structure of Hilbert space imposes strong constraints [54]. It cannot be cloned, broadcasted, or erased. It is nobody’s knowledge: it is neither epistemic nor ontological. It behaves as an independent physical entity. And what most important quantum information can be measured, processed (manipulated) stored [55, 56, 57, 1, 58, 59].
The discovery that quantum information is a new subtle resource for non-classical tasks [60, 61, 62, 63, 64, 65, 66, 67, 68, 69] revealed its operational meaning, which was of great importance for the development of quantum information theory and quantum technologies. One of the arguments in favor of the reality of quantum information is the resource criterion: If a quantum property associated with the system can be used as resource in non-classical tasks it is related to the physical reality [70].
However, the problem is that quantum information is relational in the sense it depends of the context [71, 72]. Hence, it has a “palette” of the properties within the different physical measurement contexts referred to methods of preparation (source information) and measurement contexts [73, 74].
On a formal level, composite quantum systems can be decomposed into subsystems in many different inequivalent ways. For example, if we decompose the Hilbert space $H$ of compound system S as a tensor product of smaller spaces $H=H_{1}\otimes H_{2},\ldots,\otimes H_{n}$ then the unit vectors and the states of the system can be very different, so different decompositions can have different properties. This difference is due to the fact that the symmetry group $U(d_{1})\times\ldots\times U(d_{k})$, $d_{1}=\dim(H_{i})$ is much smaller than $U(d)$, $d=\prod_{i=1}^{k}d_{i}$.The particular decomposition for the compound system represents meronomic reference frame. It may happen that in one meronomic system, separable (product) states may become entangled and vice versa. [75, 76]. It can be implemented in various possible ways, and an observer can have access to the system via control and measurement interactions associated with a corresponding experimental setup.
For quantized electromagnetic fields, the “quantum” and “classical” features of quantum information depend on the preparation of the initial states and the method of detection [77]. Namely, in contrast to classical communication where the information is encoded in classical properties of electromagnetic field (intensity or phase) quantum communication needs to send quantum information carriers such as nonclassical states of light e.g., squeezed states or Fock states that carry a well-defined number of photons. In the typical communications links, the task is an optimization over all ensembles of input quantum states used to carry information under relevant physical constraints. It is the Holevo theorem that provides a tight bound on the mutual information attainable for a given ensemble of input quantum states.
Although the notion of quantum information was known before [78], it was only the concept of transmitting quantum states intact in the papers of the teleportation [15] and Schumacher quantum coding [79] that was revolutionary, because at that time the extremely epistemic Copenhagen’s interpretation was dominant. According to it, the wave function $\psi$ (quantum state) is merely the description of the measuring and preparing apparatuses. From such a point of view, it was rather impossible to think that the message could be something else than Shannon’s classical information. However, the correspondence principle prefers interpretation according to which: the wave function $\psi$ (the quantum state) is a mathematical representation (picture) of quantum information.
Note that the above interpretation goes beyond the dilemma of the Scylla of ontology and the Charybdis of instrumentalism, and it is consistent with the well-known fact that in quantum models, the fundamental symmetry constraints are imposed on the quantum state rather than on probabilities.
To our knowledge, there no formal, commonly accepted definition of quantum information. Intuitively, one can think that the quantum information is what is carried by quantum systems in the given setup context. Nevertheless, there are fundamental “no go” broadcasting theorems [80, 81] that impose strong constraints on the processing of quantum information. More precisely, there are no completely positive trace-preserving maps that can duplicate any unknown quantum states. Perfect broadcasting is impossible because for the non-commuting states two copies (however correlated) will have more information than a single copy. For pure states, broadcasting is equivalent to cloning. The difference between cloning and broadcasting is that in cloning, one is required to obtain independent copies: In broadcasting, correlations between the copies are allowed [70].
In contrast to an “unspeakable” quantum information, its mathematical representation $\psi$ can be broadcasted. A paradigmatic example is quantum cryptography considered in [82]. Namely, the key distribution schemes show that there is a feature that is naturally ascribed to the quantum information carriers (QIC). It is clear that there must be quantum communication between Alice and Bob to achieve secret key. Indeed, Alice cannot simply send a wave function $\psi$ written on paper to Bob, as Eve could read it out without disturbing it. However, by sending QIC, Alice and Bob can achieve something impossible within the “classical” world. Another example is the direct calculation of quantum state function $f(\varrho)$ [82]. It can be viewed as a kind of quantum computing on purely quantum input. This approach is natural if the quantum state $\varrho$ is viewed as representing real quantity to be processed rather than state of our knowledge.
The above examples show that the quantum information exists physically, but it cannot be identified with its mathematical representation $\psi$. Here, a natural question arises: How does this affect the standard understanding of the Born algorithm? It is not excluded that in contrast to the classical models in which the concept of the classical information and probability are on the same footing in the sense that they can be derived from one another. At the quantum level, the concept of probability has its formal root in Gleason’s theorem, while the notion of quantum information is still not well defined. The reason is that in the quantum models of closed systems, there is no general concept of joint and conditional probability cannot be defined for the general case except for the commuting observables [78]. Indeed, in the RC model a probabilistic measure defined on phase space differs radically from the quantum probabilistic measure. In fact, the latter is largely determined by the projection structure of the von Neumann algebra, and as it follows from Gleason’s theorem, therefore it cannot be freely defined. Quantum observables are not random variables in classical probabilistic space; there are no “elements of reality.” However, for the classical measure, there is great freedom in the way it is applied from the outside to the structure of the space of phase states. It seems that the “Born’s noncontextual probability” is not an autonomic notion but depends on nonclassical algebraic properties of quantum observables.
6 Nonrelativistic quantum model: laboratory space versus configuration space
It is still remains unclear why the non-relativistic standard model already predicts the existence of stable quantum structures and EPR correlations and at the same time is consistent with RQF. Note that to protect the later from superluminal speeds, a priory postulate of local commutativity is necessary [83]. However, the RQF is not strictly limited by the speed of light [84]. In this context, the existence of the EPR correlations naturally lead to the question of their physical nature and still is subject to a wide debate [85]. In particular, Gisin suggested “that quantum correlations somehow arise from outside space-time”[86]. Clauser and Shimony had previously suggested that “Either we must completely abandon the realistic philosophy of almost all practicing physicists or fundamentally revise our concept of space-time” [87]. The second part of the above alternative is big challenge since Lorenz’s symmetry $M_{4}$ imposed by heave classical devices is an a priori measurement condition of any physical model (Bohr). On the other hand this challenge encounters an intriguing problem of incompatibility formulations of NRQ in laboratory space and configuration space, which has recently been the subject of deep consideration in the context of local realism and correlation EPR [88].
Namely, according to Born’s original formulation, the wave function $\psi_{lab}(\mathbf{r}_{lab},t)$ represents a field that propagates in lab space and correspondingly the real-valued probability density $|\psi_{lab}(\mathbf{r}_{lab},t)|^{2}$ also propagates as a classical field. Thus, Born’s modification of the interpretation of the wave function is expressed only in its statistical interpretation. This is what limits lab-space formulation the most is that it only applies to the single particle systems, and there is no strict method to extend this formulation to $n\geq 2$ particle systems. In this context, Clauser notes “This impossibility might be expected since Local Realism’s formulation and Bell’s Theorem forbids it. Correspondingly, experiments that refute Local Realism also refute a lab space formulation of quantum mechanics.”
Von Neumann introduced the abstract configuration-space formulation of quantum mechanics for the system consisting of $N$ particles, which uses a very general abstract $k$-dimensional argument space for the wave function that is called configuration space ($k=3N$), which enables a system with $k$-degrees of freedom: $\mathbf{r}_{1,conf},\ldots,\mathbf{r}_{N,conf}$, and this was generalized by incorporating $k$-quantum degree of freedom of the system: [89] $s_{1,conf},\ldots,s_{N,conf}$. Then, a configuration-space wave function can be written as: $\psi_{conf,N}(\mathbf{r}_{1,conf},\ldots,\mathbf{r}_{N,conf},s_{1,conf},\ldots,s_{N,conf})$. As Clauser notes, Born’s Lab-space formulation can only describe single-particle systems and it is essentially a form of BCHS local realism that is violated by quantum mechanics, while a completely abstract configuration space approach grasps the holistic phenomenon of entanglement, which is not a one-body phenomenon but always occurs in multipartite systems.
The inconsistency between the formulation of the laboratory space and the concept of configuration space, a.k.a Born’s argument space ambiguity shows implicitly that the fundamental dispute between Einstein and Bohr regarding the interpretation of the wave function was in fact, a dispute between Einstein, the “eventist” (“Alles ist Koinzidenz”) and Bohr, the “epistemic relationist”. Einstein believed that since measurement is subject to Lorentz $M_{4}$ symmetry, hence the classical “esse” coincides with “percipi” (measurement) in the laboratory space and this view does not contradict Born’s Lab-space statistical interpretation. In this context, it may be surprising that both model RQ and NRQ maintain the classical space-time continuum of directly observable events as an elementary background in the most fundamental models.
Whereas Bohr does not associate any reality with the wave function and treats it as intersubjective information about what can be updated in the measurement. Bohr $s$ view seems to be consistent with the configuration-space formulated wave function, as the latter does not provide a space-time description of the global wave function. Clauser concludes “if we assume, according to the Copenhagen interpretation of quantum mechanics, that abstract configuration space is inherently impenetrable, then it allows for magical action at a distance, and no objects are necessary to exist as things in the laboratory space” [88].
7 Quantum relational space hypothesis
Note, however, that if we accept the correspondence principle, we can further claim that the wave function $\psi_{conf}$ is a mathematical representation (image) of quantum information. Consequently, it must have a non-spatial nature. Unfortunately the $\psi_{conf}$ space formalism does not provide any conceptual model to understand its “inner workings”, although intuitively, it seems to hide a deeper level of reality. Despite the vast knowledge about entanglement, it is not still clear where it has its habitat and how it “coexists” with fundamental Galilean and Lorentz-Poincare symmetries. In 1999, Gisin’s team conducted a Michelson-Morely-like experiment which showed that the 2-photon interferences were still visible, independently of the relative velocity between Alice’s and Bob’s reference frames (actually, the magnitude of the velocity was constant at $100m/s$, but its orientation varied) [90] This result suggests that the von Neumann’s collapse of the quantized electromagnetic field cannot be viewed as a real causal process, and quantization of the field definitely renders it no longer describable as “real stuff in real space–time” [90, 88].
The above conclusions incline us to postulate that the fundamental physical interactions between quantum objects generate directly unobservable relations, being a source of their physical properties. In particular, we assume that every completely isolated system can be associated with an intrinsic quantum space $Q_{rel}$ of directly unobservable relations generated by unitary interactions which generate a broad class of relations in which no subsystems are required to be macroscopic directly accessible to active observer. We assume that the space of relations $Q_{rel}$ involves intrinsic quantum degrees of freedom such as spin, color, and smell, and it precedes the space of directly observable correlations, $Q_{cor}$, which is some subspace of the “classical” Minkowski $M_{4}$ space-time (Fig.1).
Figure 1: $Q_{rel}$ relation space of reversible “pre-measurements” precedes the space of correlation events $Q_{cor}$. Arrows indicate irreversible decoherence
We are not able to determine the mathematical structure of the hypothetical relational space $Q_{rel}$. The simplest way to get around this difficulty is to assume that the relational space structure is isomorphic to the Hilbert space tensorial structure. Then, we can say that quantum models provide instructions on how to model a structure of $Q_{rel}$ in terms of Hilbert space formalism. For instance, the algorithmic generation of genuine multipartite entangled states via indistinguishability of identical particles by coherent unitary deformation operations can be viewed as the kind of modeling of the $Q_{rel}$ space [91]. In particular, pure quantum state $\psi$ of the compound system encodes information about all relations (amplitude and phase) between its subsystems, which may be related to “the correlations without correlata” [9]. Obviously, the assumption on the ideal isolation of the system is realistic idealization, as the unitary interactions do not yield outcomes [92] and the system cannot “recognize” itself. It is intimately connected with the measurement problem – “Achilles’ heel” of quantum modeling of the world [28, 29, 30]. It should be noted here that for additive conserved quantities, the measurement problem is, in principle, solvable within ”Stinesprig” unitary representation based on dilated Hilbert space that includes both the degrees of freedom of the specimen and apparatus [93, 94]. In this approach, irreversible collapse does not appear from the Schrödinger equation but is induced then temporal evolution is constrained by conservation laws. In realistic scenarios, however, the observer is not able to control all the degrees of freedom environment of the system and broadcast of the results [95]. Thus, he is forced to use a formalism that includes open systems and irreversible decoherence (i.e., [96, 97]). Some authors assume [8, 98] that any physical system can play the role of an observer in a physical interaction that every interaction is a measurement regardless of the strength and time of the interaction between the systems and Bohr’s heavy basis conditions [99], which says that the apparatus is capable of realizing ideal space-time realization only if it is infinitely inert (See in this context [100, 101]).
Here, we distinguish a class of projective, strong, and instantaneous measurements that satisfy the condition of heavy basis, which removes the quantum Doppler effect. This leads to a boundary factual-like type of correlations between a ”specimen” and measuring device where the latter is a subsystem large enough that an observer can perceive it directly. It is compatible with our assumption that the relation space $Q_{rel}$ precedes subspace $Q_{cor}$ of Minkowski space $M_{4}$ of the actualizations (clicks) appearing as a result of irreversible measurements with specially prepared macro-objects (detectors) in finite regions $C$ of $M_{4}$. Then, it is reasonable to introduce measure $\mu(C)$ as the probability of a ‘click’ provided that a suitable detector operates in region C [102]. This seems consistent with the suggestion [103] that the quantum probabilistic structure is “overbuilt” over $M_{4}$, while the latter is only a “carrier” for the probabilistic measures. Note, however, that the Hilbertian “isomorphic constraint” imposed on the structure of $Q{rel}$ may be too restrictive, which may rule out the description of potential “exotic” effects such as quantum gravity.
Finally, it is worth noting, that the hypothesis of relational space confirms Gisin’s conjecture that the “quantum correlations arise from outside space-time” and that idea of quantum relational space is compatible with Reichenbach common cause principle [104].
8 Chyliński’s quantum relational continuum space
In this section, we present Chyliński’s model of a relational continuum space based on the configuration space of relative coordinates. The NRQ model ($\hbar\neq 0$) maintains the classical continuum of directly observable events $X^{\mu}=(\mathbf{X},X_{0})$, $\mu=1,2,3,4$, as an elementary background of most physical models. On the other hand, according to the uncertainty relations, the exact determination of 4-momentum $P^{\mu}=(\mathbf{P},P_{0})$, $P^{\mu}(\Delta P^{c}\to 0)$ rules out the $x$ localization of its quantum carrier in $M_{4}$ since $\Delta X^{\mu}\sim h/\Delta P^{\mu}\to\infty$. It destroys in an uncontrollable way the “quantum-potential” structure $\psi$ of the isolated system under measurement. As one knows, the direct p-measurements of the scattered particles in the asymptotic region do not disturb the structure of $\psi$ state, which is consistent with the state of isolation of the investigated system. The typical momentum measurements require measuring position and inferring momentum, which, however, do not violate the uncertainty principle. Note that there are also momentum measurements that do not rely on measuring position based on the Doppler effect, the energy-sensitive single-photon detector[105].
It is not obvious that $L$-symmetry of “heavy bases” valid micro-states when it is extrapolated to the interaction region. As we know the quantum $p$-$x$ duality establishes one-to-one correspondence between the p and x representations of a Hilbert vector $|G\rangle$ as
[…]
Note, however, that in this abstract formulation of the state $G$ of system $S_{M}$ the physical meaning of the $x$ and $p$ variables remain undetermined. According to the $M_{4}$ framework, the $p$’s denote four-momenta of the corresponding degree of freedom of $S_{M}$ represented in some reference frame $S$. However, 4-symmetry revealed the experimental role of the asymptotic language $p$ of relativistic kinematics, which was enhanced by relativistic S matric formalism. Thus, we can first measure the $S$ matrix elements (cross-sections) parametrized by Mandelsztam $M$-invariant variable $s_{k}$ determined by asymptotic 4-momenta of free particles of the initial $|i\rangle$ and final $|f\rangle$ asymptotic states of the collision process
$S_{fi}=\delta^{(4)}(P_{i}-P_{f})T_{fi}(s_{1},\ldots,s_{k}),\quad s=(P_{i}-P_{f})^{2}$ (2)
where $P_{i}$ and $P_{f}$ denote 4-momenta of the whole fully isolated system inside which the quantum-potential collision takes place. Having cross-sections, we can evaluate from (3) the $x$ representation of the corresponding structures, which is determined in the $p$ language. In particular, the S matrix formalism was used to estimate (in a non-relativistic approximation) the scattering cross-sections in collisions of electrons with atomic nuclei, based on an original Chyliński [43] model of internal quantum relation continuum $Q_{rel}$ space. To better understand the later let us briefly recall the singularity of Galilean space $G_{4}$. For this purpose, consider the simplest compound 2-body system $S_{M}=S_{1}+S_{2}$ configuration space $E_{3}(\mathbf{r}_{1})\times E_{3}(\mathbf{r}_{2})$ based on an arbitrary reference frame $S$ parametrizing $G_{4}$. The $G$-absolute time parametrizes evolution of all possible of $S_{M}$ generated by Hamiltonian
$H^{G}=\mathbf{P}_{1}^{2}/2m_{1}+\mathbf{P}_{2}^{2}/2m_{2}+V((\mathbf{r}_{2}-\mathbf{r}_{1})^{2})$ (3)
which depends on $S$ via kinetic energy term while the potential $V$ is a $G$-absolute relational shape. Using the well-known point and canonical transformations
$\mathbf{y}^{G}$ $=$ $\mathbf{r}_{2}^{G}-\mathbf{r}_{1}^{G},\mathbf{R}^{G}=a\mathbf{r}_{1}^{G}+(1-a)\mathbf{r}_{2}^{G};\;a=m_{1}/M$ (4)
$\mathbf{q}^{G}$ $=$ $a\mathbf{p}^{G}-(1-a)\mathbf{p}_{1}^{G},\quad\mathbf{P}^{G}=\mathbf{p}_{1}^{G}+\mathbf{p}_{2}^{G}$ (5)
where $\mathbf{p}^{G}$, $\mathbf{P}^{G}$ are the momenta canonically conjugate to $\mathbf{y}^{G}$, $\mathbf{R}^{G}$ respectively, we are passing to the new meronomic frame reference for the system $S_{M}$ in which the internal ($\mathbf{y}^{G}$, $\mathbf{q}^{G}$) and external ($\mathbf{R}^{G}$, $\mathbf{P}^{G}$) degree of freedoms respectively are dynamically separated. Note that the $\mathbf{y}_{G}$ and $\mathbf{q}^{G}$ are invariants of the Galilean transformation between reference frame $S_{M}$, $S^{\prime}_{M}$: $\mathbf{y}^{G}=\mathbf{y}^{G^{\prime}}$ and $\mathbf{q}^{G}=\mathbf{q}^{G^{\prime}}$, hence they measure neither the localization of $S_{M}$ nor its velocity in the $G_{4}$. In result the total operator Hamiltonian $\widehat{H}^{G}$ of $S_{M}$ of after Schrödinger quantization in the $G_{4}$ space can be written in the form:
$\widehat{H}^{G}=\hat{h}^{G}(\mathbf{y}^{G},\mathbf{q}^{G})+(\widehat{\mathbf{P}}^{G})^{2}/2M$ (6)
where that the Galilean absolute internal Hamiltonian $\hat{h}^{G}$ becomes Galilean-form-invariant. Consequently, the energy eigenstates of the $S_{M}$ are product states
$|\psi^{G}(\mathbf{R}^{G};\mathbf{y}^{G},\mathbf{q}^{G})\rangle=|\psi^{G}(\mathbf{R}^{G})\rangle\otimes|\psi^{G}(\mathbf{y}^{G},\mathbf{q}^{G})\rangle$ (7)
which are plane wave states for the center of mass and bound and unbound states of the relative motion. However, these eigenstates are entangled states of the two original subsystems of $S_{M}$
In particular, the internal states of $S_{M}$ are the solution of the Schrödinger equation
$i\hbar\frac{\partial}{\partial\tau^{G}}|\psi^{G}\rangle=\hat{h}^{G}|\psi^{G}\rangle$ (8)
where $\hat{h}^{G}$ in the internal Hamiltonian operator.
Due to the separability of the internal from the external degrees of freedom of $S_{M}$ it is convenient to introduce the internal relation continuum space the $\mathbf{Q}_{rel,4}^{G}$ composed of internal space $\mathbf{Q}_{rel,3}^{G}$ spanned by a relative coordinate $\mathbf{y}^{G}$ and the internal time $\tau^{G}$ which up to the translation constant, can be identified with the absolute Newtonian time $t^{G}$. It is not hard to see that the symmetry of (5) consists of rotations in $\mathbf{Q}_{rel,3}^{G}$ can be interpreted as the symmetry of empty space $\mathbf{Q}_{rel,4}^{G}$. Consequently all internal characteristic of $S_{M}$ in $\mathbf{Q}_{rel,3}^{G}$ are Galilean absolute. Note that $\mathbf{Q}_{rel,4}^{G}$ and $G_{4}$ deal with the same Galilean absolute intervals $r^{G}=|\mathbf{y}^{G}|$ and $\Delta r^{G}$, in addition the points $(\mathbf{y}^{G},\tau^{G})$ of $\mathbf{Q}_{rel,4}^{G}$ are determined through the point (events) of $G_{4}$. Thus the G-symmetry admits a relational absolute space $\mathbf{Q}_{rel,4}^{G}$ which is the Cartesian product $\mathbf{Q}_{rel,4}^{G}=\mathbf{Q}_{rel,3}\times T$, which operates with two absolute intervals
$r^{G}=|\mathbf{y}^{G}|\quad\textnormal{and}\quad\Delta\tau^{G}=\Delta t^{G}$ (9)
and it is isomorphic to the Galilean $G_{4}$ space-time since.
$\mathbf{y}^{G}=(\mathbf{r}_{2}^{G}-\mathbf{r}_{1}^{G})|_{\Delta t^{G}=0},\quad\Delta\tau^{G}=\Delta t^{G}$ (10)
Thus, the peculiarity of geometry $G_{4}$ comes down to the fact that the Galilean “eventism” coexists with the “relationism” of $\mathbf{Q}_{rel,4}$. The isomorphy of $\mathbf{Q}_{rel,4}$ to $G_{4}$ implies that $G_{4}$ gives room to both states and processes in the nonrelativistic limit $c\to\infty$.
8.1 Internal relation quantum space time $\mathbf{Q}_{rel,4}$
A hypothesis of the relation continuum $\mathbf{Q}_{rel,4}$ is closely connected with the isolation of physical systems, which extends to finite universal constant $c$ the absolute nature of Galilean relative coordinates and absolute Newtonian time. To extend the absoluteness of $\mathbf{Q}_{rel,4}^{G}$, let us consider first the block dual space-times $\widetilde{M}_{4}(x_{\mu})$ and $\widetilde{M}_{4}(p_{\mu})$ where $x_{\mu}=X_{\mu}^{(2)}-X_{\mu}^{(1)}$ and $p_{\mu}=P_{\mu}^{(2)}-P_{\mu}^{(1)}$ are relative coordinates and 4-momenta, respectively, whose shapes are related to each other by the unitary transformation $U$. Note that the points of the block (indivisible) space-time $\widetilde{M}_{4}(x_{\mu})$ remain completely unlocalized in measurement space-time $M_{4}$
Now consider the directly measurable observables $\widetilde{G}(p_{\mu}^{2})=\widetilde{G^{\prime}}(p_{\mu}^{\prime 2})$, assuming that their $x$ representations are well defined, i.e., the corresponding integrals converge, in the asymptotic zone $M$-form-invariant function. Then the $M_{4}-x$ representation of Hilbert vector $|G\rangle$ is given by
$\langle x|G\rangle=(2\pi\hbar)^{-4}\int_{D}d^{4}p\widetilde{G}(p_{\mu}^{2})\exp[i(p_{\mu}x_{\mu})/\hbar]$ (11)
where $\xi_{\mu}$ are two arbitrary four points (events) in $M_{4}$ and $p_{\mu}$ are relative 4-momenta of the whole fully isolated system. Here we assume that (13) is well determined under a proper contour $D$ in the $p_{0}$ – complex plane.
Definition 1.
The three-dimensional Lorentz absolute momentum space $\mathbf{Q}_{rel,3}$ spanned by the momentum $\mathbf{q}$ is defined as a space in which the auxiliary function $\tilde{F}(q^{2})$ is embedded so that, for spacelike $p_{\mu}$
$\widetilde{F}(\mathbf{q}^{2})=\widetilde{G}(p_{\mu}^{2})=p_{\mu}^{\prime 2}=\mathbf{q}^{2}\geq 0)$ (12)
Here $p_{\mu}^{2}$, $p_{\mu}^{\prime 2}$ are representations of the 4-momentum in arbitrary frames $S,S$. Note that the above equality which defines internal space $\mathbf{Q}_{rel,3}$ does not distinguish any reference frame $S$ in $M_{4}$. By definition $\widetilde{F}(\mathbf{q}^{2})$ is implicitly Lorentz-form-invariant whereas $\widetilde{G}(p_{\mu}^{2})$ is explicitly Lorentz-form-invariant. In this way, the directly measurable functions $G$ implicitly the determine $L$ absolute 3-dimensional space $Q_{rel,3}$. In accordance with the quantum canonical symmetry $\tilde{F}(q)$ is interpreted as the momentum representation in $\mathbf{Q}_{rel}$ of the Hilbert vector $F$ space $|F\rangle$, $\widetilde{F}(\mathbf{q}^{2})=\langle\mathbf{q}|F\rangle$
$\langle\mathbf{y}|F\rangle=F(\mathbf{y}^{2})=(2\pi\hbar)^{-3}\int d^{3}q\widetilde{F}(\mathbf{q}^{2})\exp[i(\mathbf{q}\mathbf{y})/\hbar]$ (13)
is implicitly Lorentz-form-invariant quantity. Finally the quantum internal relational space $\mathbf{Q}_{rel,3}$ spanned by the vector $\mathbf{y}$ is defined as the three-dimensional $\mathbf{x}$ space.
Note that the class of $|F\rangle$’s is restricted according to the convergence of the corresponding integrals. In particular interval $|\mathbf{y}|=F(\mathbf{y}^{2})=r$ in $\mathbf{Q}_{rel,3}$ has no representation in $M_{4}$ while the interval $x^{2}=G(x^{2})$ has no representation in $\mathbf{Q}_{rel,3}$. However, all typical bound states and Yukawa-type potentials have both internal and external representations in $\mathbf{Q}_{rel,4}$ and $L_{4}$.
Now, according to the quantum canonical symmetry $\mathbf{Q}$ in the Schrödinger representation $\hat{\mathbf{y}}=\mathbf{y}$, $\hat{\mathbf{q}}=-i\hbar\textnormal{grad}_{y}$ one can introduce an $M$-absolute internal Hamiltonian operator $\widehat{h}(\hat{\mathbf{y}},\hat{\mathbf{q}})$ in the configuration space $\mathbf{Q}_{rel,3}$, and the corresponding Schrödinger equation of the internal motion of an isolated system $S_{M}$ can be written in the form:
$i\hbar\frac{\partial}{\partial\tau}|\psi\rangle=\hat{h}|\psi\rangle$ (14)
where the internal Hamiltonian $\hat{h}$ is spanned by $\hat{\mathbf{y}}$ and $\hat{\mathbf{q}}$ and, consistent with the relativistic kinematics, defines the absolute evolution parameter $\tau$ which completes the space $\mathbf{Q}_{rel,4}$ into the 4-dimensional continuum of the internal space-time $S_{M}$ and hence it is the Cartesian product $\mathbf{Q}_{rel,4}=\mathbf{Q}_{rel,3}\times T$ which operates with the two absolute intervals $r=|\mathbf{y}|$, $\Delta\tau=\Delta t$, which coincide with the geometry intervals $\mathbf{Q}_{rel,4}$ (10). Thus the nonrelativistic limit of $\mathbf{Q}_{rel,3}$ is the space $\mathbf{Q}_{rel,3}^{G}$ immersed in the Galilean space-time, while the internal time $\tau$ coincides up to an arbitrary translation constant with the $G$-absolute Newtonian time.
Chyliński argued that the space $\mathbf{Q}_{rel,3}$ provides the first background of directly unobservable structures of quantum microsystems, and the $L$-symmetry becomes a limiting symmetry of $\mathbf{Q}_{rel,3}$ conditioned by infinite inertia of interacting constituents including measuring devices in the asymptotic zone of scattering. An important novelty introduced by $\mathbf{Q}_{rel,3}$ is the nonlocality of the L-form invariant form factors $G(x^{2})$, which allows correlations over spatial intervals. Moreover, Chyliński showed, that the concept of $\mathbf{Q}_{rel,4}$ space leads to potentially measurable effects in elastic, inelastic and dissociation only in the processes in which composite particles participate. Therefore, the quantum electrodynamics of point particles is left unmodified.
It is interesting that this very rough model, ignoring the quark structure of the nucleon for elastic form factor electron–nucleon collisions model predicts proton mean-square radius 0.81fm, while recent experiments provide values: 0.843fm [106], 0.833fm [107, 108] 0.8414fm [109]. These experimental discrepancies indicate that the proton radius problem still remains open.
9 Conclusion
The concept of quantum information is central to the quantum model. In particular, it is shown that “many fundamental features of quantum systems can be understood as consequences of the constraints that limited information content imposes on the systems” [46]. However, this does not translate easily into the interpretive problems of quantum theory. We try to determine the status of quantum information and its place based on two concepts. The first is the correspondence principle between the physical structure represented by the set of observer-independent data, including fundamental constants of nature, and the mathematical structure of the model that it describes. This principle allows us to interpret the wave function as a mathematical representation (image) of quantum information that is relational in nature. Consequently, we introduced the concept of relational space, in which quantum relations precede correlations, and we postulated that its mathematical structure is isomorphic to the tensorial Hilbert space structure. This interpretation becomes particularly interesting in the context of the inconsistency between Born’s interpretation of the wave function in lab space and von Neumann’s configuration space. In this approach, correlations are interpreted as limiting relations requiring decoherence and irreversibility of recorded results. Hence, this approach naturally excludes Wigner’s type paradoxes. Finally, we consider Chyliński’s quantum relational continuum space model, which leads to potentially observable effects in elastic, inelastic, and dissociation processes only for bound states.
Unfortunately, Chyliński’s articles went unnoticed, and to our knowledge, the theoretical predictions have not been tested experimentally. Even though 40 years have passed, the idea of Chyliński’s internal quantum space still seems interesting, especially in the context of the quantum correlations phenomena.
[…]
Acknowledgments The author would like to especially thank L. Newelski for his suggestion to replace the isomorphic principle with a correspondence principle that goes beyond a purely mathematical context. Special thanks are due to A. Wójcik for a panoramic, deep insight into the idea of this paper, and M. Eckstein and K. Horodecki for their deep and constructive criticism. The author would also like to thank M. Eckstein and P. Horodecki for a sequence of inspiring discussions and also M. Czachor and M. Nowakowski for their discussion, and Ł. Pankowski for help with editing the manuscript. The work was partly supported by the Foundation for Polish Science (IRAP Project, ICTQT, Contract No. MAB/2018/5, co-financed by EU within Smart Growth Operational Programme).